Grade 11Physical Chemistry

If 10^21 molecules are removed from 200mg Of Co2. How many moles will be left?

Amit
10 Years agoGrade 11

2 Answers

Himanshu
10 Years ago
No of moles of CO2 initially = weight of substance(gm)/ molecular mass
No.of moles of CO2 initially = 0.2/44 = 1/220 mol
No. of moles removed = no. of molecules/ avagardo constant
No. of moles removed = 1021/ (6 × 1023) = 1/600 mol
No. of moles left = 1/220 – 1/600 = 0.028 (approx.)
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem.
 
No of moles of CO2 initially = mass of CO2(gm)/ molecular mass
No.of moles of CO2 initially = 0.2/44 = 1/220 mol = 4.5454 x 10-3 mol
No. of moles removed = no. of molecules/ avagardo constant
No. of moles removed = 1021/ (6.022 × 1023) = 1.6606 x 10-3 mol
No. of moles left = 4.5454 x 10-3 – 1.6606 x 10-3 = 0.00288 moles (approx.)
 
Thanks and regards,
Kushagra